# Propagation of Singularities for the Damped Stochastic Klein-Gordon Equation

### v1

@@ Abstract
For the $1+1$ dimensional damped stochastic Klein-Gordon equation, we show that random singularities associated with the law of the iterated logarithm exist and propogate in the same way as the stochastic wave equation. This provides evidence for possible connections to microlocal analysis, ie. the exact regularity and singularities described in this paper should admit wavefront set type descriptions whose propagation is determined by the highest order terms of the linear operator. Despite the results being exactly the same as those of the wave equation, our proofs are significantly different than the proofs for the wave equation. Miraculously, proving our results for the critically damped equation implies them for the general equation, which significantly simplifies the problem. Even after this simplification, many important parts of the proof are significantly different than (and we think are more intuitive from the PDE viewpoint compared to) existing proofs for the wave equation.
@@ 1 Introduction
One of the ingredients for proving Theorem 1.4 is the following simultaneous upper bound for the law of the iterated logarithm. For notational convenience, we introduce a change of coordinates via a rotation by $45^{\circ}$:
@@
In general one can define $\mathop{}\!\mathbin{\Box}=b^{2}\partial_{tt}-c^{2}\partial_{xx}$, but a simple change of coordinates would allow one to produce the above results (with appropriately tilted characteristic lines) with the same calculations. We also note that all of our results can apply to the excited stochastic Klein-Gordon equation $(a<0)$, but with extra dependence on $T$ the time horizon in all of the constants in the estimates.
@@
Theorems 1.11.2 together imply that the $B^{s}_{\infty,\infty}$-Besov wavefront set [DRS22] of $u$ is empty if $s<\frac{1}{2}$ and is equal to $(\mathbb{R}_{+}\times\mathbb{R})\times(\mathbb{R}^{2}\setminus\{(0,0)\})$ if $s\geq\frac{1}{2}$. With some extra work, the same statements can be said about the more familiar $H^{s}$-Sobolev wavefront sets [Hin25, Definition 7.35, Proposition 7.36] of $u$ with the same regularity parameters. We will discuss the relevance of this in the literature review.
@@ 1.1 Literature review and motivation
$(\partial_{t}-\mathop{}\!\mathbin{\bigtriangleup}-m^{2})v(t,x)=\dot{W}(t,x)$ (9)
@@
$(\mu^{2}\partial_{tt}+\partial_{t}-\partial_{xx}-m^{2})v_{\mu}(t,x)=\dot{W}(t,x),\quad\mu\in(0,1]$ (10)
where one will see (9) as $\mu\downarrow 0$ (we refer interested readers to [CF06, CS16, CX23] for the stability of this approximation and related topics). For $\mu>0$, one can easily adopt the contents of this paper to see that Theorems 1.11.4 all apply to $v_{\mu}$ with appropriately tilted characteristic lines. The effect of tilting the characteristic lines as $\mu$ gets smaller is that for any singularity of $v_{\mu}$ at some point $(t_{0},x_{0})$, the propagated singularity of part (ii) of Theorem 1.4 is at $(t_{0}+\varepsilon,x_{0}+\frac{\varepsilon}{\mu})$ at time $t_{0}+\varepsilon$, which notably goes to infinity as $\mu\downarrow 0$. This is consistent with the PDE heuristic that hyperbolic operators have finite speed of propagation while parabolic operators have infinite speed of propagation, so $\mathop{}\!\mathbin{\Box}_{\mu}:=\mu^{2}\partial_{tt}-\partial_{xx}$ indexes a family of hyperbolic operators whose speed of propagation goes to infinity as $\mu\downarrow 0$.
@@ 3 Reduction to the critical damping case
Fix $z_{0}>0$ and assume that Theorem 1.3 holds for $u_{C}$. This implies that there exists a constant $0<K_{2}<\infty$ such that $\mathbb{P}(A_{N})=1$ for all $N\in\mathbb{N}_{+}$, where
$A_{N}=\left\{\ell_{C}(w,z_{0})\leq K_{2}\sqrt{w+z_{0}}\text{ for all $w\in[0,N]$}\right\}.$
@@
$F=\left\{(t,x)=(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}):w\in[0,N],z=z_{0}\right\}$
shows that there is an event $B_{N}$ with $\mathbb{P}(B_{N})=1$ on which (20) holds. It is clear that
$B_{N}\subset\left\{\ell_{L}(w,z_{0})=0\text{ for all $w\in[0,N]$}\right\}.$
It follows from (17) and (21) that on the event $A_{N}\cap B_{N}$, which has probability 1, we have
$\ell(w,z_{0})\leq\ell_{C}(w,z_{0})+\ell_{L}(w,z_{0})\leq K_{2}\sqrt{w+z_{0}}\quad\text{for all $w\in[0,N]$.}$
Finally, we may take $\bigcap_{N\in\mathbb{N}_{+}}(A_{N}\cap B_{N})$ as the desired event for which (6) holds.
@@
Denote by $\tilde{F}[f](\tau,\xi)$ the space-time fourier transform of an arbitrary $f(t,x)$ and $\tilde{\Gamma}(\tau,\xi):=\tilde{F}[\Gamma(\cdot,\circ)](\tau,\xi)$.
@@
Note here that $\tilde{\phi}_{t}$ is a map $\mathbb{C}\supset\mathbb{R}+i\frac{a}{2}\to\mathbb{C}$, so we need not worry about $\tilde{y}=0$. However, it is easy to show using the elementary arguments that
@@
which completes the proof of part (a).
For part (b), we write
@@
This theorem is an analogue of [DMX15, Lemma 9.3], and the original proof can be re-written here with usage of Lemmas 4.2 and 4.3 in the appropriate places with no changes in the current setting. ∎
@@
We will now show that all the integrals above are $\lesssim\varepsilon_{1}\varepsilon_{2}$ with the constant possibly depending on $T$, which will conclude the proof.
For $\tilde{I}_{12}$, 11-$\infty$ Hölder with the boundedness of the integrand and the area of $\tilde{R}_{12}$ gives the result.
@@
This proves part (i) of Theorem 1.4.
To prove part (ii), let $Z$ be an $\mathscr{F}_{t_{0}}$-measurable random variable such that

### v2

@@ Abstract
For the $1+1$ dimensional damped stochastic Klein-Gordon equation, we show that random singularities associated with the law of the iterated logarithm exist and propagate in the same way as the stochastic wave equation. This provides evidence for possible connections to microlocal analysis, i.e., the exact regularity and singularities described in this paper should admit wavefront set type descriptions whose propagation is determined by the highest order terms of the linear operator. Despite the results being exactly the same as those of the wave equation, our proofs are significantly different than the proofs for the wave equation. Miraculously, proving our results for the critically damped equation implies them for the general equation, which significantly simplifies the problem. Even after this simplification, many important parts of the proof are significantly different than (and we think are more intuitive from the PDE viewpoint compared to) existing proofs for the wave equation.
2020 Mathematics Subject Classification: 60H15, 60G17.
Keywords: stochastic partial differential equation, Klein-Gordon equation, singularity, law of the iterated logarithm
@@ 1 Introduction
One of the ingredients for proving Theorem 1.4 is the following simultaneous upper bound for the law of the iterated logarithm. For notational convenience, we introduce a change of coordinates:
@@
The statements of Theorem 1.1 and 1.2 as well as part (i) of Theorem 1.4 can be extended to statements about increments in all directions as opposed to only the characteristic directions. To not make the paper unnecessarily complicated, we chose to avoid pursuing these extensions. These extensions of Theorems 1.11.2 together imply that with probability 1, the $B^{s}_{\infty,\infty}$-Besov wavefront set [DRS22] of $u$ is empty if $s<\frac{1}{2}$ and is equal to $(\mathbb{R}_{+}\times\mathbb{R})\times(\mathbb{R}^{2}\setminus\{(0,0)\})$ if $s\geq\frac{1}{2}$. With some extra work, the same statements can be said about the more familiar $H^{s}$-Sobolev wavefront sets [Hin25, Definition 7.35, Proposition 7.36] of $u$ for the same ranges of $s$. We will discuss the relevance of this in the literature review.
@@
In general one can define $\mathop{}\!\mathbin{\Box}=b^{2}\partial_{tt}-c^{2}\partial_{xx}$, but a simple change of coordinates would allow one to produce the above results (with appropriately tilted characteristic lines) with the same calculations. In particular, what one would see is that singularities where the increments are observed in one characteristic direction would be propagated forward in time in the other characteristic direction. If one studies the proof of Theorem 1.3 and the propagation part of 1.4, one will discover that the proofs would fail if one tries to repeat it for any other pair of unit vectors that are not the characteristic pair. This is also predictable from the known behavior of hyperbolic equations, and we will discuss more of its relevance in the literature review. We also note that all of our results can apply to the excited stochastic Klein-Gordon equation $(a<0)$, but with extra dependence on $T$ the time horizon in all of the constants in the estimates.
@@ 1.1 Literature review and motivation
$(\partial_{t}-\mathop{}\!\mathbin{\bigtriangleup})v(t,x)=\dot{W}(t,x)$ (9)
@@
$(\mu^{2}\partial_{tt}+\partial_{t}-\partial_{xx})v_{\mu}(t,x)=\dot{W}(t,x),\quad\mu\in(0,1]$ (10)
where one will see (9) as $\mu\downarrow 0$ (we refer interested readers to [CF06, CS16, CX23] for the stability of this approximation and related topics). For $\mu>0$, one can easily adopt the contents of this paper to see that Theorems 1.11.4 all apply to $v_{\mu}$ with appropriately tilted characteristic lines. The effect of tilting the characteristic lines as $\mu$ gets smaller is that for any singularity of $v_{\mu}$ at some point $(t_{0},x_{0})$, the propagated singularity of part (ii) of Theorem 1.4 is at $(t_{0}+\varepsilon,x_{0}+\frac{\varepsilon}{\mu})$ at time $t_{0}+\varepsilon$, which notably goes to infinity as $\mu\downarrow 0$. This is consistent with the PDE heuristic that hyperbolic operators have finite speed of propagation while parabolic operators have infinite speed of propagation, so $\mathcal{L}_{\mu}:=\mu^{2}\partial_{tt}+\partial_{t}-\partial_{xx}$ indexes a family of hyperbolic operators whose speed of propagation goes to infinity as $\mu\downarrow 0$.
@@ 3 Reduction to the critical damping case
Fix $N>0$, $0<a_{1}<a_{2}$ and assume that Theorem 1.3 holds for $u_{C}$. This implies that there exists a constant $0<K_{2}<\infty$ such that $\mathbb{P}(A_{z_{0}})=1$ for all $z_{0}\in[a_{1},a_{2}]$, where
$A_{z_{0}}=\left\{\ell_{C}(w,z_{0})\leq K_{2}\text{ for all $w\in[0,N]$}\right\}.$
@@
$F=\left\{(t,x)=(\tfrac{w+z}{\sqrt{2}},\tfrac{-w+z}{\sqrt{2}}):w\in[0,N],z\in[a_{1},a_{2}]\right\}$
shows that there is an event $B$ with $\mathbb{P}(B)=1$ on which (20) holds. It is clear that for any $z_{0}\in[a_{1},a_{2}]$,
$B\subset\left\{\ell_{L}(w,z_{0})=0\text{ for all $w\in[0,N]$}\right\}.$
It follows from (17) and (21) that on the event $A_{z_{0}}\cap B$, which has probability 1, we have
$\ell(w,z_{0})\leq\ell_{C}(w,z_{0})+\ell_{L}(w,z_{0})\leq K_{2}\quad\text{for all $w\in[0,N]$.}$
@@
Thanks to the reduction in Section 3, we assume in the remainder of the paper that $m^{2}=a^{2}/2$. Denote by $\tilde{F}[f](\tau,\xi)$ the space-time fourier transform of an arbitrary $f(t,x)$ and $\tilde{\Gamma}(\tau,\xi):=\tilde{F}[\Gamma(\cdot,\circ)](\tau,\xi)$.
@@
Note here that $\tilde{\phi}_{t}$ is a map $\mathbb{C}\supset\mathbb{R}+i\frac{a}{2}\to\mathbb{C}$, so we need not worry about $\tilde{y}=0$. However, it is easy to show using elementary arguments that
@@
which completes the proof of part (a). For part (b), we write
@@
This theorem is an analogue of [DMX15, Lemma 9.3], and the original proof can be re-written here with usage of Lemmas 4.2 and 4.3 (with $\alpha=1/2$) in the appropriate places with no changes in the current setting. ∎
@@
We will now show that all the integrals above are $\lesssim\varepsilon_{1}\varepsilon_{2}$ with the constant possibly depending on $T$, which will conclude the proof. For $\tilde{I}_{12}$, 11-$\infty$ Hölder with the boundedness of the integrand and the area of $\tilde{R}_{12}$ gives the result.
@@
This proves part (i) of Theorem 1.4. To prove part (ii), let $Z$ be an $\mathscr{F}_{t_{0}}$-measurable random variable such that
@@
Acknowledgment. The second author is supported in part by a research startup fund of the Chinese University of Hong Kong, Shenzhen and the Shenzhen Peacock fund 2025TC0013.